Can someone help me solve this for X:
4x/x-4 +1 = 3x/x-4
A)x= 2
B)x= -2
C)x= -1/2
D)x= 1/2
Please show me your work so that I can learn how to do this.

Answer by stanbon(57963)

4x/x-4 +1 = 3x/x-4
Multiply thru by (x-4) to get:
4x + (x-4)= 3x
5x-4=3x
2x=4
x=2
Cheers,
Stan H.